# Large language model psychometrics: A systematic review of evaluation, validation, and enhancement

### v2

@@ 1 Introduction
The advent of large language models (LLMs) represents a transformative breakthrough in AI. These systems exhibit general-purpose capabilities spanning diverse domains [Bubeck et al., 2023], with particular proficiency in natural language understanding and generation [Demszky et al., 2023, Grossmann et al., 2023, Ziems et al., 2024, Gu et al., 2024]. They are rapidly integrated into critical societal infrastructure, ranging from consumer-facing applications like chatbots [OpenAI, 2025] and search engines [Wang et al., 2024c] to high-stakes domains such as healthcare [Singhal et al., 2023], education [Milano et al., 2023], and scientific discovery [Romera-Paredes et al., 2024, Ye et al., 2024]. Their increasing dominance has exposed a fundamental, pressing scientific challenge: how can we rigorously evaluate these AI systems that transcend traditional benchmarks of biological or algorithmic intelligence?
@@
These challenges intersect with humanity’s century-old quest to quantify the complex, intangible human psychology [Pasquali, 2009]. Psychometrics emerged from this timeless pursuit as the scientific study of psychological measurement. It bridges the abstract and the empirical by transforming human traits into quantifiable data, enabling better understanding, prediction, and decision-making in education, business, healthcare, governance, and beyond [Rust and Golombok, 2014]. The intersection of LLMs and psychometrics forms a methodological crucible, forging novel paradigms to better decode and improve machine minds.
@@
This intersection sets the stage for a new research frontier. We define LLM Psychometrics as the interdisciplinary field dedicated to evaluating, understanding, and enhancing LLMs through the application and integration of psychometric instruments, theories, and principles. This field seeks to quantify, interpret, manipulate, and improve the complex, human-like attributes and behaviors exhibited by LLMs, encompassing both personality constructs (e.g., personality, values, morality, and attitudes) and cognitive constructs (e.g., heuristics and biases, social interaction abilities, psycholinguistic abilities, and learning and cognitive capacities). Research in LLM Psychometrics applies, extends, and innovates upon scientific methods of psychological measurement for LLMs. Based on the psychometric principles, related research systematically examines the measurement results to ensure scientific rigor. By measuring and elucidating psychological constructs in LLMs, LLM Psychometrics further informs strategies for their targeted enhancement.
@@
Recent research in LLM Psychometrics pioneers in addressing the LLM evaluation crisis. Some studies introduce dynamic and construct-oriented evaluation frameworks that move beyond static, task-specific benchmarks [Zhu et al., 2024a, Hagendorff, 2023]. In parallel, novel methodologies are developed to measure non-cognitive and emergent constructs [Ren et al., 2024, Huang et al., 2023d, Pellert et al., 2024]. Self-adaptive evaluation techniques now allow for the extrapolation of item difficulty and tailoring assessments to model performance [Jiang et al., 2024a, Lalor et al., 2024, Polo et al., 2024]. Drawing from the methodological framework of psychometric validation, research improves the reliability and validity of evaluation protocols [Ye et al., 2025a]. Human-centered evaluation drives aligning model behavior with human values [Wang et al., 2024f, Yao et al., 2025a]. In addition, the scope of evaluation expands to agentic and multimodal systems, further broadening the methodological landscape [Li et al., 2024b, Huang et al., 2024c].
@@
The field of LLM Psychometrics has seen significant growth, evidenced by the proliferation of related research papers. These studies, however, address a variety of psychological constructs, employ diverse methodologies, and utilize distinct validation techniques. The interdisciplinary nature of this domain has attracted contributions from a broad spectrum of academic communities. Despite this diversity, there is a lack of cohesion among researchers, leading to a fragmentation of insights, particularly between studies focusing on disparate constructs. Consequently, there is a pressing need for a systematic review to synthesize these efforts and facilitate a more integrated understanding of the field.
@@
To bridge the gap, we systematically review LLM Psychometrics across evaluation, validation, and enhancement. The evaluation framework encompasses three core dimensions: the target construct (what to measure), the measurement method (how to measure), and the validation of results (how well do we measure). The psychometric insights not only inform evaluation but also guide the development and refinement of LLMs (how to improve). Accordingly, this review is structured as follows (Fig. 1). § 2 provides an overview of the preliminaries and methodological foundations to facilitate subsequent discussions. § 3 thoroughly contrasts psychometrics and traditional AI benchmarking and reviews how psychometric principles can underpin and reinvent LLM benchmarks. The core measurement framework is detailed in § 4, § 5, and § 6. § 4 delves into the psychological constructs evaluated in LLMs, elucidating the theories employed and summarizing key evaluation findings. In § 5, we scrutinize the psychometric evaluation methodologies applied to LLMs, followed by § 6, which examines the psychometric validation of evaluation results. Beyond evaluation, § 7 introduces strategies for enhancing LLMs through psychometric insights. § 8 discusses emerging trends, challenges, and future directions. Finally, § 9 highlights key ethical considerations, while § 10 concludes the paper.
@@
We consider studies to be out of scope if they do not employ psychometric approaches, adhere to psychometric principles, or if they focus solely on scalar performance metrics rather than characterizing behaviors. Readers interested in conventional LLM benchmarking are encouraged to consult surveys on LLM evaluation [Guo et al., 2023b, Chang et al., 2024]. Several related reviews focus on the evaluation of specific constructs in LLMs, such as personality [Wen et al., 2024b, Dong et al., 2025], attitudes and values [Ma et al., 2024a], cultural awareness [Pawar et al., 2024, Adilazuarda et al., 2024], and theory of mind [Sarıtaş et al., 2025, Dong et al., 2025]. Hagendorff [2023], Hagendorff et al. [2024] introduce the notion of machine psychology and review emergent LLM abilities; however, they do not provide comprehensive coverage of the related research, nor do they elaborate on LLM personality constructs, psychometric validation, or enhancement. This paper provides the first systematic review of LLM Psychometrics.
@@ 2 Preliminary and methodological foundations
We aim for our survey to be self-contained and accessible to a broad, cross-disciplinary audience. To this end, this section presents the preliminary and methodological foundations that underpin the subsequent discussions.
@@ 2.1 Large language models
A key emergent capability of LLMs is in-context learning, which allows models to adapt to new tasks or patterns by conditioning on examples or instructions provided within the input context $x_{<t}$, without modifying model parameters. This property can influence LLM performance in psychometric evaluation. For instance, prompting models to reason step-by-step (e.g., Chain-of-Thought prompting [Wei et al., 2022]) can enhance performance on reasoning tasks, while instructing them to role-play may modulate their exhibited personalities and values.
@@ 2.2 Psychometrics
A psychological test is a set of items that are designed to measure characteristics of human beings that pertain to behavior [Kaplan and Saccuzzo, 2001]. Behavior measured by tests can be overt (observable actions) or covert (internal thoughts or feelings). Tests may assess past, current, or even predict future behavior. Interpretation of test scores depends on their context within a distribution. Scales are used to relate raw scores to defined distributions, aiding interpretation. In addition, psychological tests can measure traits—enduring tendencies like shyness or determination—and states, which reflect temporary conditions of individuals.
Psychological testing measures individual differences in various constructs, which are abstract psychological attributes or dimensions that help explain and predict behavior. Two primary categories of such constructs are personality constructs and cognitive constructs [Kaplan and Saccuzzo, 2001]. Personality tests focus on an individual’s tendencies and dispositions. These tests measure typical behavior, such as preferences or tendencies to react in certain ways. Cognitive tests evaluate speed, accuracy, or both, with higher scores reflecting better performance.
Two fundamental principles underpin psychometrics: reliability and validity [Raykov and Marcoulides, 2011]. Reliability ensures accuracy, dependability, consistency, or repeatability of the test results. Reliable test results are stable across time, contexts, and raters. Validity confirms the meaningfulness and usefulness of the test results. A valid measure captures the intended construct. Validity is multifaceted; for example, predictive validity might correlate test scores with job performance, while construct validity ensures alignment with theoretical models, such as the Big Five personality traits [Goldberg, 2013].
Other principles include standardization, which provides context to raw scores by comparing individual results to a representative sample, or norm group. Additionally, equivalence and fairness are crucial principles that tests must adhere to. Test bias occurs when test items unintentionally advantage or disadvantage subgroups. Modern psychometrics employs advanced statistical models to identify and revise biased items, ensuring assessments measure the intended construct rather than extraneous factors [Rust and Golombok, 2014].
@@ 2.3 Psychometric evaluation of AI before the era of LLMs
The idea of applying psychometrics to AI originated in the early decades of AI [Pellert et al., 2024]. Evans [1964] pioneered work in this area by creating a heuristic program that could solve parts of intelligence tests. Subsequent efforts similarly focused on designing AI systems for cognitive tests [Newell, 1973], with the goal of creating systems capable of handling human tasks. This was conceptually aligned with the development of static, task-centric benchmarks in modern AI research [Hendrycks et al., 2020, Chen et al., 2021, Liang et al., 2022, Srivastava et al., 2022, Lee et al., 2024e]. However, criticisms emerged regarding the absence of "hot cognition" in AI, prompting Simon [1963] to propose incorporating emotional aspects into models. By the early 2000s, the concept of "psychometric AI" was explicitly articulated as the pursuit of systems capable of excelling on all established, validated tests of intelligence and mental ability. These included not only conventional IQ tests but also assessments of artistic and literary creativity, mechanical ability, and beyond [Bringsjord and Schimanski, 2003, Pellert et al., 2024]. It was not until the advent of LLMs that the versatility envisioned for "psychometric AI" began to materialize.
@@ 3.1 Fundamental differences between psychometrics and AI benchmarking
Benchmarking AI systems superficially resembles psychometrics, particularly Classical Test Theory (CTT) [Crocker and Algina, 1986], as both compile test items to evaluate cognitive capabilities and average the resulting scores. However, closer examination reveals that AI benchmarks differ significantly from modern psychometric approaches [Wang et al., 2023a, Federiakin, 2025]. We outline these key differences in Table 1.
@@
Psychometrics centers on understanding the psychological constructs and ensuring that tests accurately measure the intended constructs. Grounded in a causal measurement philosophy, this field posits that observed test responses arise from latent psychological constructs [Markus and Borsboom, 2013, Federiakin, 2025]. These constructs may encompass both abilities (e.g., reasoning skills) and personalities (e.g., conscientiousness). The causal framework necessitates rigorous construct definition, requiring traits to be precisely delineated through iterative theory-building and empirical validation. Psychometric test development follows methodical protocols, often structured by frameworks such as Evidence-Centered Design (ECD) [Mislevy et al., 2003]. ECD emphasizes ensuring congruence between test items and theoretical models of the construct, thereby supporting robust inferences about latent traits.
Central to this approach is the prioritization of item quality over quantity. Psychometricians conduct rigorous item analyses to balance precision with practicality, as administering an excessive number of items to participants is often impractical. Advanced statistical models, such as Item Response Theory (IRT) [Embretson and Reise, 2013] and Factor Analysis [Loehlin, 2004], are adopted to estimate latent traits, analyze item performance, and assess model fit. These models require relatively large sample sizes (the number of human participants) to yield stable parameter estimates, as they must disentangle individual differences from measurement error to accurately infer latent traits.
The test results are analyzed to ensure the reliability, validity, predictive power, and explanatory power of the test. Specifically, a well-designed test should: 1) consistently and accurately measure the intended construct, 2) predict performance across a diverse range of related tasks and real-world outcomes, and 3) provide explanatory insight into the observed data. For example, psychometric models often reveal that individual differences across a broad range of cognitive tasks can be captured and explained by a relatively small set of underlying cognitive abilities [Cattell and Horn, 1978].
@@
In contrast, AI benchmarking is driven by pragmatic goals: evaluating and ranking models based on task performance. Unlike psychometrics, validity is not the primary concern. Instead, benchmarks typically emphasize breadth, scalability, and—especially in the era of foundation models—difficulty. This approach reflects a representativist philosophy, where it is assumed that an extensive set of benchmark items collectively captures all relevant aspects of the abilities demanded by the target task [Federiakin, 2025]. However, constructs like reasoning or knowledge are often ambiguously defined and encompass infinitely many aspects. Benchmarks implicitly operationalize these constructs through ad hoc task selection.
The development of AI benchmarks is usually less labor-intensive, especially when compared with psychometrics on a per-item basis. Test items and their corresponding ground truths are typically drawn from existing datasets, expert curation, or crowdsourced contributions. While this process enables scalability, it risks conflating superficial task performance with deeper cognitive capacities. For instance, a benchmark may assess mathematical reasoning through arithmetic problems without verifying whether models rely on pattern recognition versus symbolic logic [Ahn et al., 2024]. Additionally, LLM benchmarking commonly employs straightforward metrics, such as average accuracy, eschewing the sophisticated latent variable models of psychometrics. This simplicity allows benchmarks to evaluate single models efficiently, bypassing the need for population samples. However, it also limits the depth of insights that can be gleaned from model performance [Federiakin, 2025].
Reliability and stability in benchmarking are primarily achieved through scaling up the test. However, ensuring the quality of each individual item becomes impractical due to the test scale and the rapid pace of model development. For instance, while psychometrics emphasizes the discriminative power of each item, some benchmarks, though initially challenging, are quickly outpaced by continuous model improvements [McIntosh et al., 2024]. Conversely, certain emerging benchmarks are currently too difficult to yield meaningful comparisons [Phan et al., 2025]. Benchmark results are often limited to the specific target task, offering limited generalizability or predictive power across other tasks or real-world applications. These results also pose significant challenges for conducting in-depth, multi-faceted analyses of model capabilities [Wang et al., 2023a].

### v3

@@ 1 Introduction
The advent of large language models (LLMs) represents a transformative breakthrough in AI. These systems exhibit general-purpose capabilities spanning diverse domains [bubeck2023sparks], with particular proficiency in natural language understanding and generation [demszky2023using, grossmann2023ai, ziems2024can, gu2024llm-as-a-judge]. They are rapidly integrated into critical societal infrastructure, ranging from consumer-facing applications like chatbots [chatgpt] and search engines [wang2024large_search_model] to high-stakes domains such as healthcare [singhal2023large], education [milano2023large], and scientific discovery [romera2024mathematical, ye2024reevo]. Their increasing dominance has exposed a fundamental, pressing scientific challenge: how can we rigorously evaluate these AI systems that transcend traditional benchmarks of biological or algorithmic intelligence?
@@
These challenges intersect with humanity’s century-old quest to quantify the complex, intangible human psychology [pasquali2009psychometrics]. Psychometrics emerged from this timeless pursuit as the scientific study of psychological measurement. It bridges the abstract and the empirical by transforming human traits into quantifiable data, enabling better understanding, prediction, and decision-making in education, business, healthcare, governance, and beyond [rust2014modern_psychometrics].
@@
This intersection sets the stage for a new research frontier that we term LLM Psychometrics. This interdisciplinary field is dedicated to evaluating, understanding, and enhancing LLMs through the application and integration of psychometric instruments, theories, and principles. Crucially, LLM Psychometrics treats LLMs themselves as the measured subjects, rather than instruments for measuring humans. The field seeks to quantify, interpret, manipulate, and improve the human-like attributes and behaviors exhibited by LLMs, spanning non-cognitive constructs (e.g., personality, values, morality, attitudes) and cognitive constructs (e.g., heuristics and biases, social interaction abilities, psycholinguistic abilities, learning and reasoning capacities). Grounded in psychometric principles, research in LLM Psychometrics examines measurement properties (e.g., reliability, validity, generalizability, and measurement invariance) and uses these insights to inform targeted model improvement.
@@
Recent research in LLM Psychometrics pioneers in addressing the LLM evaluation crisis. Some studies introduce dynamic and construct-oriented evaluation frameworks that move beyond static, task-specific benchmarks [zhu2024dynamic, hagendorff2023machine]. In parallel, novel methodologies are developed to measure non-cognitive constructs [ren2024valuebench, huang2023psychobench, pellert2024ai_psychometrics]. Self-adaptive evaluation techniques now allow for the extrapolation of item difficulty and tailoring assessments to model performance [jiang2024raising, lalor2024item, polo2024tinybenchmarks]. Drawing from the methodological framework of psychometric validation, research improves the reliability and validity of evaluation protocols [ye2025gpv]. Human-centered evaluation drives aligning model behavior with human values [wang2024incharacter, yao2025value_compass_leaderboard]. In addition, the scope of evaluation expands to agentic and multimodal systems, further broadening the methodological landscape [li2024quantifying, huang2024designing].
Notably, we view LLM Psychometrics as a methodological borrowing focused on behavioral manifestation. We do not posit that LLMs possess subjective experience, sentience, or developmental histories. Rather, our focus is on how psychological constructs manifest in their outputs, as well as the fidelity, consistency, and controllability of these behavioral manifestations. Accordingly, when we refer to a construct in the context of LLMs throughout this review, we mean a synthetic behavioral manifestation: systematic response patterns measurable via psychometric frameworks, rather than a claim about an underlying genuine psychological state. Given the increasing integration of LLMs into human-facing roles, measuring and controlling these manifested behaviors holds immense practical value for user experience, safety, and alignment, regardless of the models’ internal ontological status. However, realizing this practical value is non-trivial. The very act of applying anthropocentric tests to non-human entities raises unresolved questions about reliability and validity. A fundamental debate continues regarding whether these instruments capture meaningful latent patterns in LLMs or merely reflect sophisticated statistical mimicry. Furthermore, the sensitivity of LLMs to prompts, decoding strategies, and context strains core psychometric assumptions such as trait stability.
@@
Amidst the debates, the field of LLM Psychometrics has seen significant growth, evidenced by the proliferation of related research papers. These studies, however, often operate in silos, addressing isolated psychological constructs, employing diverse methodologies, and utilizing distinct validation techniques. The interdisciplinary nature of this domain has attracted contributions from a broad spectrum of academic communities, yet there remains a lack of cohesion among researchers. This fragmentation of insights, particularly between studies focusing on disparate constructs, underscores a pressing need for a systematic review to synthesize these efforts and facilitate a more integrated, scientifically robust understanding of the field.
@@
To bridge the gap, we systematically review LLM Psychometrics across evaluation, validation, and enhancement. The evaluation framework encompasses three core dimensions: the target construct (what to measure), the measurement method (how to measure), and the validation of results (how well do we measure). The psychometric insights not only inform evaluation but also guide the development and refinement of LLMs (how to improve). Accordingly, this review is structured as follows (Fig.˜1). §˜2 provides an overview of the preliminaries and methodological foundations to facilitate subsequent discussions. §˜3 thoroughly contrasts psychometrics and traditional AI benchmarking and reviews how psychometric principles can underpin and reinvent LLM benchmarks. The core measurement framework is detailed in §˜4, §˜5, and §˜6. §˜4 delves into the psychological constructs evaluated in LLMs, elucidating the theories employed and summarizing key evaluation findings. In §˜5, we scrutinize the psychometric evaluation methodologies applied to LLMs, followed by §˜6, which examines the psychometric validation of evaluation results. Beyond evaluation, §˜7 introduces strategies for enhancing LLMs through psychometric insights. §˜8 discusses emerging trends, challenges, and future directions. Finally, §˜9 highlights key ethical considerations, while §˜10 concludes the paper.
@@
This review treats LLMs as subjects of psychometric tests. We distinguish this from using LLMs as tools to assess or profile human respondents, which is out of scope; see ye2025psychometrics for a review of that adjacent field. We also exclude evaluation studies if they do not employ psychometric approaches, adhere to psychometric principles, or if they focus solely on scalar performance metrics rather than characterizing latent behavioral tendencies. The operational distinction between psychometric research and conventional AI benchmarking, which underpins these inclusion criteria, is elaborated in §˜3.1. Studies that merely administer scales and report scores meet the minimal inclusion threshold but offer limited scientific value. Rigorous LLM Psychometrics additionally evaluates measurement properties, such as reliability across prompt variants and construct validity via factor analysis. Readers interested in conventional LLM benchmarking are encouraged to consult surveys on LLM evaluation [guo2023survey_llm_eval, chang2024survey_llm_eval]. Several related reviews focus on the evaluation of specific constructs in LLMs, such as personality [wen2024llm_personality_survey, dong2025humanizing], attitudes and values [ma2024aov_survey], cultural awareness [pawar2024culture_survey, adilazuarda2024culture_survey], and theory of mind [saritacs2025tom_survey, dong2025humanizing]. In addition, fairness and related psychometric concepts (e.g., differential item functioning, measurement invariance) are primarily discussed in the context of using LLMs as tools for human psychometrics. We refer readers interested in AI fairness to dedicated surveys [mehrabi2021survey, gallegos2024bias_fairness_survey]. hagendorff2023machine, hagendorff2024machine introduce the notion of machine psychology and discuss the emergent LLM abilities; however, they do not provide comprehensive coverage of the related research, nor do they elaborate on LLM personality constructs, psychometric validation, or enhancement. This paper provides the first systematic review of LLM Psychometrics.
@@ 2 Preliminary and methodological foundations
We aim for our review to be self-contained and accessible to a broad, cross-disciplinary audience. To this end, this section presents the preliminary and methodological foundations that underpin the subsequent discussions.
@@ 2.1 Large language models
A key emergent capability of LLMs is in-context learning, which allows models to adapt to new tasks or patterns by conditioning on examples or instructions provided within the input context $x_{<t}$, without modifying model parameters. This property can influence LLM performance in psychometric evaluation. For instance, prompting models to reason step-by-step (e.g., Chain-of-Thought prompting [wei2022chain]) can enhance performance on reasoning tasks, while instructing them to role-play may modulate their exhibited personalities and values.
@@ 2.2 Psychometrics
A psychological test is a set of items that are designed to measure characteristics of human beings that pertain to behavior [kaplan2001psychological_testing_book]. Behavior measured by tests can be overt (observable actions) or covert (internal thoughts or feelings). Tests may assess past, current, or even predict future behavior. Interpretation of test scores depends on their context within a distribution. Scales are used to relate raw scores to defined distributions, aiding interpretation. In addition, psychological tests can measure traits—enduring tendencies like shyness or determination—and states, which reflect temporary conditions of individuals.
Psychological testing measures individual differences in various constructs, which are abstract psychological attributes or dimensions that help explain and predict behavior. Two primary categories of such constructs are personality constructs and cognitive constructs [kaplan2001psychological_testing_book]. Personality tests focus on an individual’s tendencies and dispositions. These tests measure typical behavior, such as preferences or tendencies to react in certain ways. Cognitive tests evaluate speed, accuracy, or both, with higher scores reflecting better performance.
Two fundamental principles underpin psychometrics: reliability and validity [raykov2011introduction_psychometric_theory]. Reliability ensures accuracy, dependability, consistency, or repeatability of the test results. Reliable test results are stable across time, contexts, and raters. Validity confirms the meaningfulness and usefulness of the test results. A valid measure captures the intended construct. Validity is multifaceted; for example, predictive validity might correlate test scores with job performance, while construct validity ensures alignment with theoretical models, such as the Big Five personality traits [goldberg2013big_five].
Other principles include standardization, which provides context to raw scores by comparing individual results to a representative sample, or norm group. Additionally, measurement invariance (or equivalence) and fairness are crucial principles that tests must adhere to. Measurement invariance ensures that the same construct is measured in the same way across different groups. Violations of this often stem from test bias, where items (referred to as showing differential item functioning or DIF) unintentionally advantage or disadvantage subgroups. Modern psychometrics employs advanced statistical models to identify and revise such biased items, ensuring assessments measure the intended construct rather than extraneous factors [rust2014modern_psychometrics]. In LLM Psychometrics, a particularly fundamental form of this concern is construct equivalence: whether a construct such as personality or values means the same thing when applied to LLMs as it does for humans.
@@ 2.3 Psychometric evaluation of AI before the era of LLMs
The idea of applying psychometrics to AI originated in the early decades of AI [pellert2024ai_psychometrics]. evans1964heuristic pioneered work in this area by creating a heuristic program that could solve parts of intelligence tests. Subsequent efforts similarly focused on designing AI systems for cognitive tests [newell1973], with the goal of creating systems capable of handling human tasks. This was conceptually aligned with the development of static, task-centric benchmarks in modern AI research [hendrycks2020mmlu, chen2021human_eval, liang2022helm, srivastava2022bigbench, lee2024vhelm]. However, criticisms emerged regarding the absence of "hot cognition" in AI, prompting simon1963 to propose incorporating emotional aspects into models. By the early 2000s, the concept of "psychometric AI" was explicitly articulated as the pursuit of systems capable of excelling on all established, validated tests of intelligence and mental ability. These included not only conventional IQ tests but also assessments of artistic and literary creativity, mechanical ability, and beyond [bringsjord2003psychometric_ai, pellert2024ai_psychometrics]. It was not until the advent of LLMs that the versatility envisioned for "psychometric AI" began to materialize.
@@ 3.1 Fundamental differences between psychometrics and AI benchmarking
Benchmarking AI systems superficially resembles psychometrics, particularly Classical Test Theory (CTT) [crocker1986introduction], as both compile test items to evaluate cognitive capabilities and average the resulting scores. However, closer examination reveals that AI benchmarks differ significantly from modern psychometric approaches [wang2023evaluating_general_purpose_ai, federiakin2025improving]. We outline these key differences in Table˜1.
@@
Psychometrics centers on understanding the psychological constructs and ensuring that tests accurately measure the intended constructs. Grounded in a causal measurement philosophy, this field posits that observed test responses arise from latent psychological constructs [markus2013frontiers, federiakin2025improving]. These constructs may encompass both abilities (e.g., reasoning skills) and personalities (e.g., conscientiousness). The causal framework necessitates rigorous construct definition, requiring traits to be precisely delineated through iterative theory-building and empirical validation. Psychometric test development follows methodical protocols, often structured by frameworks such as Evidence-Centered Design (ECD) [mislevy2003ecd]. ECD emphasizes ensuring congruence between test items and theoretical models of the construct, thereby supporting robust inferences about latent traits.
Central to this approach is the prioritization of item quality over quantity. Psychometricians conduct rigorous item analyses to balance precision with practicality, as administering an excessive number of items to participants is often impractical. Advanced statistical models, such as Item Response Theory (IRT) [embretson2013item_response_theory] and Factor Analysis [loehlin2004factor_analysis], are adopted to estimate latent traits, analyze item performance, and assess model fit. These models require relatively large sample sizes (the number of human participants) to yield stable parameter estimates, as they must disentangle individual differences from measurement error to accurately infer latent traits.
The test results are analyzed to ensure the reliability, validity, predictive power, and explanatory power of the test. Specifically, a well-designed test should: 1) consistently and accurately measure the intended construct, 2) predict performance across a diverse range of related tasks and real-world outcomes, and 3) provide explanatory insight into the observed data. For example, psychometric models often reveal that individual differences across a broad range of cognitive tasks can be captured and explained by a relatively small set of underlying cognitive abilities [cattell1978check].
@@
In contrast, AI benchmarking is driven by pragmatic goals: evaluating and ranking models based on task performance. Unlike psychometrics, validity is not the primary concern. Instead, benchmarks typically emphasize breadth, scalability, and—especially in the era of foundation models—difficulty. This approach reflects a representativist philosophy, where it is assumed that an extensive set of benchmark items collectively captures all relevant aspects of the abilities demanded by the target task [federiakin2025improving]. However, constructs like reasoning or knowledge are often ambiguously defined and encompass infinitely many aspects. Benchmarks implicitly operationalize these constructs through ad hoc task selection.
The development of AI benchmarks is usually less labor-intensive, especially when compared with psychometrics on a per-item basis. Test items and their corresponding ground truths are typically drawn from existing datasets, expert curation, or crowdsourced contributions. While this process enables scalability, it risks conflating superficial task performance with deeper cognitive capacities. For instance, a benchmark may assess mathematical reasoning through arithmetic problems without verifying whether models rely on pattern recognition versus symbolic logic [ahn2024llm_math_survey]. Additionally, LLM benchmarking commonly employs straightforward metrics, such as average accuracy, eschewing the sophisticated latent variable models of psychometrics. This simplicity allows benchmarks to evaluate single models efficiently, bypassing the need for population samples. However, it also limits the depth of insights that can be gleaned from model performance [federiakin2025improving].
Reliability and stability in benchmarking are primarily achieved through scaling up the test. However, ensuring the quality of each individual item becomes impractical due to the test scale and the rapid pace of model development. For instance, while psychometrics emphasizes the discriminative power of each item, some benchmarks, though initially challenging, are quickly outpaced by continuous model improvements [mcintosh2024inadequacies_of_llm_bench]. Conversely, certain emerging benchmarks are currently too difficult to yield meaningful comparisons [phan2025humanity]. Benchmark results are often limited to the specific target task, offering limited generalizability or predictive power across other tasks or real-world applications. These results also pose significant challenges for conducting in-depth, multi-faceted analyses of model capabilities [wang2023evaluating_general_purpose_ai].